OpenStudy (anonymous):

The product of two consecutive integers is 210. Find the integers. (Consecutive means following, like 11, 12)

OpenStudy (anonymous):

a(a+1)=210 a^2+a-210=0 (a+15)(a-14)=0 a=-15 or 14 so the integers are -15,-14 or 14,15